# Agentar-Fin-R1: Enhancing Financial Intelligence through Domain Expertise, Training Efficiency, and Advanced Reasoning

### v1

@@ Abstract
Large Language Models (LLMs) demonstrate tremendous potential in the financial domain, yet existing models often fall short in scenarios demanding robust reasoning capabilities, stringent trustworthiness requirements, and efficient adaptation to task-specific needs. We introduce the Agentar-Fin-R1 series of financial large language models (8B and 32B parameters), specifically engineered based on the Qwen3 foundation model to enhance reasoning capabilities, reliability, and domain specialization for financial applications. Our optimization approach integrates a high-quality, systematic financial task taxonomy with a comprehensive multi-layered trustworthiness assurance framework. This framework encompasses high-quality trustworthy knowledge engineering, multi-agent trustworthy data synthesis, and rigorous data validation governance. Through label-guided automated difficulty-aware optimization, tow-stage learning processes, and detailed attribution systems, we achieve substantial improvements in training efficiency. Our models undergo comprehensive evaluation on mainstream financial benchmarks including FinEva, FinEval, and FinanceIQ, as well as general reasoning datasets such as MATH-500 and GPQA. To thoroughly assess real-world deployment capabilities, we innovatively propose the Finova evaluation benchmark, which focuses on agent-level financial reasoning and compliance verification. Experimental results demonstrate that Agentar-Fin-R1 not only achieves state-of-the-art performance on financial tasks but also exhibits exceptional general reasoning capabilities, validating its effectiveness as a trustworthy solution for high-stakes financial applications.
@@ 1 Introduction
Figure 1: Benchmark performance of Agentar-Fin-R1
Large Language Models (LLMs) have demonstrated remarkable capabilities in complex reasoning tasks, with recent advances in reasoning-optimized models such as OpenAI’s o1 series [17], QwQ [19], DeepSeek-R1 [10], and Qwen3 [27] achieving significant breakthroughs in mathematics, programming, and logical inference. However, the direct deployment of these general-purpose models in financial applications reveals critical limitations: insufficient domain-specific financial knowledge integration, leading to poor performance on finance-related tasks; susceptibility to hallucinations that violate the stringent safety and compliance requirements essential in financial environments.
Existing financial LLMs can be broadly categorized into two types. Non-reasoning financial models such as Baichuan [29], XuanYuan [6], and PIXIU [24] incorporate domain-specific financial knowledge but lack sophisticated analytical and reasoning capabilities required for complex financial decision-making scenarios involving multi-step analysis, risk assessment, and strategic planning. Reasoning-enhanced financial models including XuanYuan-FinX1-Preview [7], Fino1 [18], Fin-R1 [15], and Dianjin-R1 [31] attempt to integrate advanced reasoning mechanisms but still exhibit significant limitations: insufficient reasoning capabilities for handling complex financial scenarios that require deep analytical thinking; lack of scenario-specific reasoning adaptation, failing to align reasoning processes with the unique demands of financial contexts such as real-time market dynamics, regulatory compliance constraints, and risk tolerance considerations.
@@
Professional Taxonomy-Guided Framework: We construct a fine-grained financial task taxonomy that decomposes the financial domain into precisely defined categories, serving as an active guidance framework throughout the entire development pipeline. This taxonomy not only guides data processing and training workflows but also enables systematic task-oriented optimization, ensuring comprehensive coverage of financial reasoning scenarios and providing professional support for model training.
@@
  • A taxonomy-guided methodology for developing trustworthy and efficient financial LLMs that systematically addresses data fragmentation, reasoning transparency, and scenario generalization challenges.
@@
Our experimental results demonstrate that Agentar-Fin-R1 achieves superior performance across financial benchmarks (Fineva, FinEval, FinanceIQ and Finova) while maintaining competitive results on general reasoning tasks (MATH, GPQA), validating the effectiveness of domain-specialized optimization without catastrophic forgetting.
@@ 2.2 Label System
  • Scene Dimension: Encompasses diverse real-world financial scenarios, including Banking, Securities, Insurance, Trusts, and Mutual Funds. Each scenario embodies a distinct application context with specialized requirements, enabling the model to adapt to the operational nuances and domain-specific characteristics of each financial sector.
  • Task Attribute Dimension: Characterizes the functional nature of tasks, encompassing Named Entity Recognition (NER), Intent Classification, Slot Filling, Entity Disambiguation, and Consultation-style Question Answering. These task attributes precisely define how input data should be interpreted and processed, providing explicit guidance for model behavior during instruction learning and response generation phases.
@@
Here, $c_{i}\in\mathcal{C}$ represents a content category (e.g., Banking, Insurance), and $a_{i}\in\mathcal{A}$ represents a task attribute (e.g., Entity Disambiguation, Slot Filling). This structured formulation facilitates fine-grained task decomposition and enables targeted data generation that is precisely aligned with downstream application requirements.
It is crucial to recognize that the Scene and Task Attribute dimensions exhibit non-orthogonal characteristics: not all task attributes are equally applicable across every scene. For instance, Slot Filling demonstrates high relevance in Banking customer service interactions but exhibits limited applicability in Trusts, where document analysis and compliance reasoning predominate. This phenomenon results in a sparsely populated cross-product space, which more authentically reflects the natural distribution and practical constraints of real-world financial tasks.
@@ 2.3.2 Synthesis: Trusted Multi-Agent Generation of Reasoning Triplets
Figure 3: Overview of the dual-track data synthesis pipeline, incorporating both knowledge-driven task agents and self-evolution mechanisms for generating verifiable reasoning triplets.
@@
This track leverages a curated financial knowledge base and a domain-specific task taxonomy to drive the generation of high-quality, verifiable (query, thinking, answer) triplets.
@@
We define a structured taxonomy of financial tasks as $\mathcal{T}=\{t_{1},t_{2},\ldots,t_{m}\}$, where each label $t_{j}\in\mathcal{T}$ represents a distinct task category (e.g., fraud detection, portfolio analysis, regulatory compliance). For each task label $t_{j}$, a dedicated generation agent $A_{t_{j}}$ is instantiated.
@@ 2.3.3 Verification and Checking: Rigorous Multi-Modal Validation
An independent verification model $M_{\text{verify}}$ validates the logical correctness of answers through systematic query analysis and reasoning process examination:
@@
The verification model generates intermediate reasoning steps to determine whether the given answer can be logically derived from the question context and available domain knowledge, ensuring adherence to financial reasoning principles.
@@
Stratified Sampling Annotation
We perform stratified random sampling on the generated data to ensure representative coverage across task types, complexity levels, and domain subcategories. The sampling ratio is carefully calibrated to balance annotation cost with statistical significance and coverage requirements. Experienced financial domain experts will conduct comprehensive multi-dimensional assessment of sampled data instances.
@@
The rating model outputs multi-dimensional quality scores encompassing accuracy, relevance, fluency, and domain-specific appropriateness metrics, providing granular quality insights.
@@ 3.1 Weighted Training Framework
Training financial large language models (LLMs) involves addressing the inherent heterogeneity and complexity of financial tasks, which exhibit varying levels of difficulty and domain-specific requirements. Traditional training methods, such as Supervised Fine-Tuning (SFT) and Reinforcement Learning (RL), treat all training samples uniformly, without accounting for the fact that some tasks are significantly more challenging than others. Consequently, models may overfit to simpler, more frequently encountered tasks while underperforming on complex tasks that are crucial for real-world financial decision-making and risk assessment.
@@
For Reinforcement Learning (RL) training phases, we modify the preference-based objective function to incorporate difficulty weights in a theoretically principled manner:
$\mathcal{L}_{\text{RL}}=-\mathbb{E}_{(x,y_{w},y_{l})}\left[\tilde{w}_{t}\cdot\log\sigma(r(x,y_{w})-r(x,y_{l}))\right]$ (17)
where $y_{w}$ and $y_{l}$ represent the "winning" and "losing" responses respectively in the preference pair, $r(\cdot)$ denotes the learned reward function, and $\sigma(\cdot)$ is the sigmoid activation function. By weighting the entire loss term rather than individual reward components, we maintain the stability and theoretical properties of the learning process while systematically emphasizing improvement on difficult tasks.
@@ 3.3 Attribution Loop
      $\Delta_{t}=\max(0,P_{\text{target}}-p_{t})$
@@
      $e_{t}=\frac{\max\left(0,\Delta p_{t}^{(k)}\right)}{\Delta d_{t}^{(k)}+\varepsilon}$
@@
      $\pi_{t}=\Delta_{t}\cdot e_{t}\cdot\exp(-\lambda d_{t})$
@@
      $\Delta d_{t}=\frac{\pi_{t}}{\sum_{t^{\prime}\in\mathcal{T}}\pi_{t^{\prime}}}\cdot B_{k}$
@@ 4.1 Benchmark
  • MATH[12]: A benchmark designed to assess a model’s ability to solve mathematical problems that require multi-step reasoning.
  • GPQA-diamond[20]: A general-purpose question-answering benchmark that tests a model’s ability to comprehend and reason through diverse, non-financial tasks.
@@ 4.2 Training detail
We train Agentar-Fin-R1-8B and Agentar-Fin-R1-32B based on Qwen3-8B-Instruct and Qwen3-32B-Instruct respectively. The training process consists of two stages: initial SFT, and then GRPO and SFT refinement. For the 8B model, we use 16 NVIDIA A100 GPUs, while the 32B model is trained on 64 A100 GPUs. All training uses bf16 precision with a sequence length of 16K and appropriate gradient accumulation steps to ensure training stability. In addition to the synthetically generated data produced through our data synthesis framework outlined in Section 2, the dataset encompasses financial reasoning data accumulated from our internal repositories[30], general-purpose training data[21], datasets from Llama-Nemotron[2] and openthoughts[9].
@@ 4.3 Baseline
  • General models without explicit reasoning: GPT-4o[16], Qwen2.5-14B-Instruct[26], Qwen2.5-72B-Instruct[26], and DeepSeek-V3[14].
  • General models with reasoning capabilities: GPT-o1[17], Qwen3-8B[27], Qwen3-32B[27], Qwen-QwQ-32B[19], and DeepSeek-R1[10].
@@ 4.4 Main Results
Model Params Financial Fin. Common Com. All
@@
Table 2: Performance comparison in accuracy across different test sets. Scores in bold indicate the best results. Agentar-Fin-R1-32B achieves state-of-the-art performance across all financial benchmarks.
@@
A critical observation is that our domain-specialized models exhibit remarkable capability preservation in general-purpose tasks, with Agentar-Fin-R1-32B attaining a 93.80 score on MATH and 68.18 on GPQA—performance levels that match or exceed those of general-purpose reasoning models with comparable parameter counts. Notably, we observed substantial improvements on GPQA compared to the base model, demonstrating that domain specialization can actually enhance performance across diverse reasoning tasks. This empirical evidence validates our hypothesis that targeted domain optimization can be achieved without compromising general cognitive capabilities, and in some cases, may even strengthen them.
@@
Table 3: Performance comparison across different capabilities. Agent Capabilities: Intent Detection(Financial Intent Detection), Slotting(Financial Slot Recognition), Tool(Financial Tool Planning), Expression(Financial Expression Generation). Complex Reasoning: Combined score of Financial Mathematics and Code Understanding and Financial Complex Reasoning. Safety & Compliance: Combined score of Safety and Compliance. Scores in bold indicate the best results.
@@
In the dimension of agent capabilities, our models demonstrate remarkable superiority in all evaluated agent capabilities. Notably, in the financial expression generation task, Agentar-Fin-R1-32B achieves an outstanding score of 69.00, significantly surpassing all competing models. This task evaluates the model’s ability to integrate complex information and generate accurate, contextually relevant expressions in financial contexts. This is particularly important as it correlates directly with hallucination suppression, a critical requirement in practical applications. The Agentar-Fin-R1 models’ ability to generate reliable, coherent, and precise financial expressions indicates their exceptional accuracy and reliability, making them highly suitable for real-world financial decision-making tasks.
In the realm of complex reasoning, which combines financial mathematics, code understanding, and intricate financial problem-solving, Agentar-Fin-R1-32B leads the way with a score of 55.64. This positions our model as the top performer in handling sophisticated financial reasoning tasks, outperforming both general-purpose models and other financial-specific models. The ability to efficiently process and solve complex reasoning tasks is vital for applications such as financial analysis, forecasting, and decision support, areas where Agentar-Fin-R1 excels due to its combination of financial expertise and advanced reasoning capabilities.
Equally significant is our model’s performance in safety and compliance tasks, where Agentar-Fin-R1-32B achieves the highest score of 87.00, far exceeding the performance of all other models. Financial systems are subject to stringent regulatory standards, and Agentar-Fin-R1 demonstrates an exceptional capacity to comply with these regulations while maintaining safety in its operations. The Agentar-Fin-R1-8B model also excels in this domain with a score of 80.50, showcasing its trustworthiness when handling sensitive financial data. These results validate our model’s application in regulated financial environments, ensuring both safety and compliance, which are paramount in any financial AI system.

### v2

@@ Abstract
Large Language Models (LLMs) exhibit considerable promise in financial applications; however, prevailing models frequently demonstrate limitations when confronted with scenarios that necessitate sophisticated reasoning capabilities, stringent trustworthiness criteria, and efficient adaptation to domain-specific requirements. We introduce the Agentar-Fin-R1 series of financial large language models (8B and 32B parameters), specifically engineered based on the Qwen3 foundation model to enhance reasoning capabilities, reliability, and domain specialization for financial applications. Our optimization approach integrates a high-quality, systematic financial task label system with a comprehensive multi-layered trustworthiness assurance framework. This framework encompasses high-quality trustworthy knowledge engineering, multi-agent trustworthy data synthesis, and rigorous data validation governance. Through label-guided automated difficulty-aware optimization, tow-stage training pipeline, and dynamic attribution systems, we achieve substantial improvements in training efficiency. Our models undergo comprehensive evaluation on mainstream financial benchmarks including Fineva, FinEval, and FinanceIQ, as well as general reasoning datasets such as MATH-500 and GPQA-diamond. To thoroughly assess real-world deployment capabilities, we innovatively propose the Finova evaluation benchmark, which focuses on agent-level financial reasoning and compliance verification. Experimental results demonstrate that Agentar-Fin-R1 not only achieves state-of-the-art performance on financial tasks but also exhibits exceptional general reasoning capabilities, validating its effectiveness as a trustworthy solution for high-stakes financial applications. The Finova bench is available at https://github.com/antgroup/Finova.
@@ 1 Introduction
Figure 1: Agentar-Fin-R1-32B performance on financial benchmarks. Results on general reasoning benchmark (MATH-500, GPQA-diamond) are detailed in following sections.
Large Language Models (LLMs) have demonstrated remarkable capabilities in complex reasoning tasks, with recent advances in reasoning-optimized models such as OpenAI’s o1 series [18], QwQ [20], DeepSeek-R1 [11],Seed-Thinking-v1.5[22] and Qwen3 [29] achieving significant breakthroughs in mathematics, programming, and logical inference. However, the direct deployment of these general-purpose models in financial applications reveals critical limitations: insufficient domain-specific financial knowledge integration, leading to poor performance on finance-related tasks; susceptibility to hallucinations that violate the stringent safety and compliance requirements essential in financial environments.
Existing financial LLMs can be broadly categorized into two types. Non-reasoning financial models such as Baichuan [31], DISC-FinLLM[4], XuanYuan [7], and PIXIU [26] incorporate domain-specific financial knowledge but lack sophisticated analytical and reasoning capabilities required for complex financial decision-making scenarios involving multi-step analysis, risk assessment, and strategic planning. Reasoning-enhanced financial models including XuanYuan-FinX1-Preview [8], Fino1 [19], Fin-R1 [16], and Dianjin-R1 [33] attempt to integrate advanced reasoning mechanisms but still exhibit significant limitations: insufficient reasoning capabilities for handling complex financial scenarios that require deep analytical thinking; lack of scenario-specific reasoning adaptation, failing to align reasoning processes with the unique demands of financial contexts such as market dynamics, regulatory compliance constraints, and risk tolerance considerations.
@@
Professional Label-Guided Framework: We construct a fine-grained financial task label system that decomposes the financial domain into precisely defined categories, serving as an active guidance framework throughout the entire development pipeline. This label system not only guides data processing and training workflows but also enables systematic task-oriented optimization, ensuring comprehensive coverage of financial reasoning scenarios and providing professional support for model training.
@@
  • A label-guided methodology for developing trustworthy and efficient financial LLMs that systematically addresses data fragmentation, reasoning transparency, and scenario generalization challenges.
@@
Our experimental results demonstrate that Agentar-Fin-R1 achieves superior performance across financial benchmarks (Fineva, FinEval, FinanceIQ and Finova) while maintaining competitive results on general reasoning tasks (MATH-500, GPQA-diamond), validating the effectiveness of domain-specialized optimization without catastrophic forgetting.
@@ 2.2 Label System
  • Scene Dimension: Encompasses diverse real-world financial scenarios, such as Banking, Securities, Insurance, Trusts, and Mutual Funds. Each scenario embodies a distinct application context with specialized requirements, enabling the model to adapt to the operational nuances and domain-specific characteristics of each financial sector.
  • Task Type Dimension: Defines the specific types of tasks to be performed, such as Named Entity Recognition (NER), Intent Classification, Slot Filling, Entity Disambiguation, and Consultation-style Question Answering. These task types specify the exact operations the model should execute on input data, providing clear directives for what actions to take during instruction learning and response generation phases.
@@
Here, $c_{i}\in\mathcal{C}$ represents a scene category (e.g., Banking, Insurance), and $a_{i}\in\mathcal{A}$ represents a task attribute (e.g., Entity Disambiguation, Slot Filling). This structured formulation facilitates fine-grained task decomposition and enables targeted data generation that is precisely aligned with downstream application requirements.
It is crucial to recognize that the Scene and Task Attribute dimensions exhibit non-orthogonal characteristics: not all task attributes are equally applicable across every scene. This phenomenon results in a sparsely populated cross-product space, which more authentically reflects the natural distribution and practical constraints of real-world financial tasks.
@@ 2.3.2 Synthesis: Trusted Multi-Agent Generation of Reasoning Triplets
Figure 3: Overview of the dual-track data synthesis pipeline, incorporating both task-oriented knowledge-guided generation and self-evolution mechanisms for generating verifiable reasoning triplets.
@@
This track leverages a curated financial knowledge base and a domain-specific task label system to drive the generation of high-quality, verifiable (query, thinking, answer) triplets.
@@
We define a structured label system of financial tasks as $\mathcal{T}=\{t_{1},t_{2},\ldots,t_{m}\}$, where each label $t_{j}\in\mathcal{T}$ represents a distinct task category (e.g., fraud detection, portfolio analysis, regulatory compliance). For each task label $t_{j}$, a dedicated generation agent $A_{t_{j}}$ is instantiated.
@@ 2.3.3 Verification and Checking: Rigorous Multi-Modal Validation
An independent third-party model validates the logical correctness of answers through a prompt-based approach, analyzing queries and reasoning processes to determine answer validity:
@@
We perform stratified random sampling of the generated data to ensure representative coverage of task types, complexity levels, and domain subcategories. The sampling ratio is carefully calibrated to balance annotation cost with statistical significance and coverage requirements. Experienced financial domain experts will conduct comprehensive multidimensional assessment of sampled data instances.
@@ 3.1 Weighted Training Framework
Training financial large language models (LLMs) involves addressing the inherent heterogeneity and complexity of financial tasks, which exhibit varying levels of difficulty and domain-specific requirements. Traditional training methods treat all training samples uniformly, without accounting for the fact that some tasks are significantly more challenging than others. Consequently, models may overfit to simpler, more frequently encountered tasks while underperforming on complex tasks that are crucial for real-world financial decision-making and risk assessment.
@@
For Reinforcement Learning (RL) training phases, we modify the preference-based objective function to incorporate difficulty weights in a theoretically principled manner.
@@ 3.3 Attribution Loop
      $\Delta_{t}=\max(0,P_{\text{target}}-p_{t})$ (18)
@@
      $e_{t}=\frac{\max\left(0,\Delta p_{t}^{(k)}\right)}{\Delta d_{t}^{(k)}+\varepsilon}$ (19)
@@
      $\pi_{t}=\Delta_{t}\cdot e_{t}\cdot\exp(-\lambda d_{t})$ (20)
@@
      $\Delta d_{t}=\frac{\pi_{t}}{\sum_{t^{\prime}\in\mathcal{T}}\pi_{t^{\prime}}}\cdot B_{k}$ (22)
@@ 4.1 Benchmark
  • MATH[13]: A benchmark designed to assess a model’s ability to solve mathematical problems that require multi-step reasoning. We used the MATH-500 subset for evaluation.
  • GPQA[21]: A general-purpose question-answering benchmark that tests a model’s ability to comprehend and reason through diverse, non-financial tasks. We used the GPQA-diamond subset for evaluation.
@@ 4.2 Training detail
We train Agentar-Fin-R1-8B and Agentar-Fin-R1-32B based on Qwen3-8B-Instruct and Qwen3-32B-Instruct respectively. The training process consists of two stages: initial SFT, and then GRPO and SFT refinement. For the 8B model, we use 16 NVIDIA A100 GPUs, while the 32B model is trained on 64 A100 GPUs. All training uses bf16 precision with a sequence length of 16K and appropriate gradient accumulation steps to ensure training stability. Beyond the synthetically generated data derived from our data synthesis framework detailed in Section 2, our dataset incorporates financial reasoning data from our proprietary Agentar-DeepFinance-300K[32], general-purpose training corpora[23], as well as datasets sourced from Llama-Nemotron[2] and openthoughts[10].
@@ 4.3 Baseline
  • General models without explicit reasoning: GPT-4o[17](Version 2024-08-06), Qwen2.5-14B-Instruct[28], Qwen2.5-72B-Instruct[28], and DeepSeek-V3[15](Version 2025-03-24).
  • General models with reasoning capabilities: GPT-o1[18](Version 2024-12-17), Qwen3-8B[29], Qwen3-32B[29], Qwen-QwQ-32B[20], and DeepSeek-R1[11](Version 2025-05-28).
@@ 4.4 Main Results
Model Params Financial Financial General General Overall
@@
Table 2: Performance comparison in accuracy across financial benchmarks (Fineva, FinEval, FinanceIQ, Finova) and general reasoning tasks (MATH: MATH-500, GPQA: GPQA-diamond). Scores in bold indicate the best results. Results include individual benchmark performance and averaged scores for financial tasks (Financial Avg.), general reasoning (General Avg.), and overall performance (Overall Avg.). Agentar-Fin-R1-32B achieves state-of-the-art performance across all financial benchmarks, as well as competitive results on general reasoning.
Figure 5: Performance comparison between Agentar-Fin-R1 and Qwen3 models (8B and 32B variants) on financial benchmarks (Fineva, FinEval, FinanceIQ, Finova) and general reasoning benchmarks (MATH: MATH-500, GPQA: GPQA-diamond). Our proposed models show improvements across both domain-specific and general reasoning tasks.
@@
A critical observation is that our domain-specialized models exhibit remarkable capability preservation in general-purpose tasks, with Agentar-Fin-R1-32B attaining a 93.80 score on MATH-500 and 68.18 on GPQA-diamond—performance levels that match or exceed those of general-purpose reasoning models with comparable parameter counts. Notably, we observed substantial improvements on GPQA-diamond compared to the base model. This empirical evidence validates our hypothesis that targeted domain optimization can be achieved without compromising general cognitive capabilities, and in some cases, may even strengthen them.
@@
Table 3: Performance comparison on Finova. Agent Capabilities: Intent(Financial Intent Detection), Slotting(Financial Slot Recognition), Tool(Financial Tool Planning), Expression(Financial Expression Generation). Complex Reasoning: Combined score of Financial Mathematics and Code Understanding and Financial Complex Reasoning. Safety & Compliance: Combined score of Safety and Compliance. Scores in bold indicate the best results.
@@
In the dimension of agent capabilities, our models demonstrate remarkable superiority in all evaluated agent capabilities. Notably, in the financial expression generation task, Agentar-Fin-R1-32B achieves an outstanding score of 69.00, significantly surpassing all competing models. This task evaluates the model’s ability to integrate complex information, contextually relevant expressions in financial contexts. This is particularly important as it correlates directly with hallucination suppression, a critical requirement in practical applications. The ability of the Agentar-Fin-R1 models to generate precise, coherent and reliable financial expressions indicates their exceptional accuracy and reliability, making them highly suitable for real-world financial decision-making tasks.
In the realm of complex reasoning, which combines financial mathematics, code understanding, and intricate financial problem solving, Agentar-Fin-R1-32B leads the way with a score of 55.64. This positions our model as the best performing model in handling sophisticated financial reasoning tasks, outperforming both general-purpose models and other financial-specific models. The ability to efficiently process and solve complex reasoning tasks is vital for applications such as financial analysis, forecasting, and decision support, areas where Agentar-Fin-R1 excels due to its combination of financial expertise and advanced reasoning capabilities.
Equally significant is our model’s performance in safety and compliance tasks, where Agentar-Fin-R1-32B achieves the highest score of 87.00, far exceeding the performance of all other models. Financial systems are subject to stringent regulatory standards, and Agentar-Fin-R1 demonstrates an exceptional capacity to comply with these regulations while maintaining safety in its operations. The Agentar-Fin-R1-8B model also excels in this domain with a score of 80.50, showcasing its trustworthiness when handling sensitive financial data. These results validate the application of our model in regulated financial environments, ensuring both safety and compliance, which are paramount in any financial AI system.